# Einstein metrics on homogeneous superspaces

### v1

@@ 4.2 Compact real forms
To construct a flag supermanifold $M=(G_{0},\mathfrak{g})/(K_{0},\mathfrak{k})$ from a given circled Dynkin diagram, we first specify the compact real form $\mathfrak{g}$. In the following, we recall Dynkin diagrams, star operations and compact real forms for the complex Lie superalgebras of types $A$ and $C$.
@@
where $[E_{ij}]=0$ if $i,j\leq m$ or $i,j>m$, and $[E_{ij}]=1$ otherwise. Let $\mathfrak{h}^{\mathbb{C}}$ be the Cartan subalgebra spanned by all diagonal matrices, and denote by $\{\varepsilon_{1},\ldots,\varepsilon_{m+n}\}$ the $(\mathfrak{h}^{\mathbb{C}})^{\vee}$-basis where $\varepsilon_{i}(E_{jj})=\delta_{ij}$ for all $i,j$. For convenience, we introduce $\delta_{\mu}:=\varepsilon_{m+\mu}$ for each $\mu=1,\ldots,n$.
Let $\mathfrak{sl}(m|n)^{\mathbb{C}}:=\{X\in\mathfrak{gl}(m|n)^{\mathbb{C}}\mid\operatorname{Str}(X)=0\}$, and let $\mathfrak{b}^{\mathbb{C}}$ be the Borel subsuperalgebra spanned by all upper-triangular matrices. Then, the set of positive roots partitions as $\Delta^{+}=\Delta_{\bar{0}}^{+}\cup\Delta_{\bar{1}}^{+}$, where
$\Delta_{\bar{0}}^{+}=\{\varepsilon_{i}-\varepsilon_{j},\delta_{\mu}-\delta_{\nu}\mid 1\leq i<j\leq m;1\leq\mu<\nu\leq n\},\qquad\Delta_{\bar{1}}^{+}=\{\varepsilon_{i}-\delta_{\mu}\mid 1\leq i\leq m;1\leq\mu\leq n\},$
@@
$\Pi=\{\alpha_{i}=\varepsilon_{i}-\varepsilon_{i+1},\alpha_{m}=\varepsilon_{m}-\delta_{1},\alpha_{m+\mu}=\delta_{\mu}-\delta_{\mu+1}\mid 1\leq i\leq m-1;1\leq\mu\leq n-1\}$
@@
The highest root is $\varepsilon_{1}-\delta_{n}=\alpha_{1}+\cdots+\alpha_{m+n-1}$, while the Weyl vector follows from
@@
$\{\imath E_{ii}\mid i=1,\ldots,m+n\}\cup\{A_{ij},B_{ij}\mid 1\leq i<j\leq m\}\cup\{A_{ij},B_{ij}\mid m+1\leq i<j\leq m+n\}$
$\cup\,\{\sqrt{\imath}A_{ij},\sqrt{\imath}B_{ij}\mid 1\leq i\leq m;\,m+1\leq j\leq m+n\},$
@@
$\Delta$ ={±$=\big{\{}\!\pm\!(\varepsilon_{1}-\varepsilon_{2}),\pm(\varepsilon_{2}-\varepsilon_{3}),\pm(\varepsilon_{1}-\varepsilon_{3}),\pm(\delta_{1}-\delta_{2})\big{\}}\cup\big{\{}\!\pm\!(\varepsilon_{i}-\delta_{1}),\pm(\varepsilon_{i}-\delta_{2})\mid i=1,2,3\big{\}},$
$\Delta_{K}$ ={±$=\big{\{}\!\pm\!(\varepsilon_{1}-\varepsilon_{2}),\pm(\delta_{1}-\delta_{2})\big{\}}\cup\big{\{}\!\pm\!(\varepsilon_{3}-\delta_{1}),\pm(\varepsilon_{3}-\delta_{2})\big{\}},$
@@
$\Delta_{M}=\Delta-\Delta_{K}=\big{\{}\!\pm\!(\varepsilon_{i}-\varepsilon_{3}),\pm(\varepsilon_{i}-\delta_{1}),\pm(\varepsilon_{i}-\delta_{2})\mid i=1,2\big{\}},$
@@
We extend the analysis above to general $\operatorname{SU}(m|n)$ with a single circled node in the Dynkin diagram in Section 4.3, and with two circled nodes in Section 4.4. A similar analysis is carried out for $\operatorname{SOSp}(2|2n)$ in Section 4.5, although for a single circled node only.
@@
$\mathfrak{osp}(2|2n)^{\mathbb{C}}:=\big{\{}X\in\mathfrak{gl}(2|2n)^{\mathbb{C}}\mid\omega(Xu,v)+(-1)^{[X][u]}\omega(u,Xv)=0\text{ for all }u,v\in V\big{\}}.$
@@
$\mathfrak{h}^{\mathbb{C}}=\big{\{}\mathrm{diag}(a,-a,f_{1},\dots,f_{n},-f_{1},\dots,-f_{n})\mid a,f_{1},\ldots,f_{n}\in\mathbb{C}\big{\}}.$ (4.10)
@@
$\Delta_{\bar{0}}^{+}=\{\delta_{i}\pm\delta_{j}\mid 1\leq i<j\leq n\}\cup\{2\delta_{j}\mid 1\leq j\leq n\},\qquad\Delta_{\bar{1}}^{+}=\{\varepsilon\pm\delta_{j}\mid 1\leq j\leq n\}.$
@@
Case $\alpha_{1}$$\alpha_{p}$$\alpha_{q}$$\alpha_{m}$$\alpha_{m+1}$$\alpha_{m+n-1}$ase 2:$\alpha_{1}$$\alpha_{p}$$\alpha_{m}$$\alpha_{m+1}$$\alpha_{q}$$\alpha_{m+n-1}$ase 3:$\alpha_{1}$$\alpha_{m-1}$$\alpha_{m}$$\alpha_{p}$$\alpha_{q}$$\alpha_{m+n-1}$
@@ 4.4.1 Isotropy summands
$\mathfrak{m}_{k_{1},k_{2}}^{\mathbb{C}}:=\bigoplus_{\alpha\in\Delta_{k_{1},k_{2}}}\mathfrak{g}_{\alpha}^{\mathbb{C}},\qquad\Delta_{k_{1},k_{2}}:=\Big{\{}\sum_{i=1}^{m+n-1}c_{i}\alpha_{i}\in\Delta_{M}\mid c_{p}=k_{1},c_{q}=k_{2}\Big{\}},\qquad(k_{1},k_{2})\in I_{\mathfrak{m}^{\mathbb{C}}},$
@@
so the coefficient of $\alpha_{p}$ is $1$, while the coefficient of $\alpha_{q}$ is 00. This implies that $\mathfrak{m}^{{}^{\prime}\mathbb{C}}_{1,0}\subseteq\mathfrak{m}^{\mathbb{C}}_{1,0}$. Similarly, it follows that $\mathfrak{m}^{{}^{\prime}\mathbb{C}}_{k_{1},k_{2}}\subseteq\mathfrak{m}^{\mathbb{C}}_{k_{1},k_{2}}$ for all $(k_{1},k_{2})\in I_{\mathfrak{m}^{\mathbb{C}}}$, so
@@
As in (2.2), $\mathfrak{m}_{\bar{0}}$ is spanned by $(-1)$-eigenvectors of the star operation restricted to $\mathfrak{m}_{\bar{0}}^{\mathbb{C}}$. Using Proposition 4.2, the star operation on $\mathfrak{m}_{\bar{0}}^{\mathbb{C}}$ splits into restrictions on $(\mathfrak{m}_{-1,0}^{\mathbb{C}}\oplus\mathfrak{m}_{1,0}^{\mathbb{C}})_{\bar{0}}$, $(\mathfrak{m}_{0,-1}^{\mathbb{C}}\oplus\mathfrak{m}_{0,1}^{\mathbb{C}})_{\bar{0}}$ and $(\mathfrak{m}_{-1,-1}^{\mathbb{C}}\oplus\mathfrak{m}_{1,1}^{\mathbb{C}})_{\bar{0}}$, respectively. For each pair $(k_{1},k_{2})\in\{(1,0),(0,1),(1,1)\}$, the real $(-1)$-eigenspace of the star operation on $(\mathfrak{m}_{-k_{1},-k_{2}}^{\mathbb{C}}\oplus\mathfrak{m}_{k_{1},k_{2}}^{\mathbb{C}})_{\bar{0}}$ is spanned by $A_{\alpha}$ and $B_{\alpha}$, $\alpha\in(\Delta_{k_{1},k_{2}}^{+})_{\bar{0}}$, so
@@
$\{A_{\alpha},B_{\alpha}\mid\alpha\in\Delta_{k_{1},k_{2}}\cap\Delta_{\bar{0}}^{+}\}\cup\{\sqrt{\imath}A_{\alpha},\sqrt{\imath}B_{\alpha}\mid\alpha\in\Delta_{k_{1},k_{2}}\cap\Delta_{\bar{1}}^{+}\}$
@@ 4.4.2 Structure constants and Ricci curvature
Case 1: $p<q\leq m$ $m-n-p$ $m-n+p-q$ $q$
Case 2: $p<m<q$ $m-n-p$ $-m-n+p+q$ $2m-q$
Case 3: $m\leq p<q$ $-m-n+p$ $m-n-p+q$ $2m-q$
@@
If $[123]\neq 0$, then $d_{1},d_{2},d_{3}$ are nonzero and the desired expression for $r_{i}$ follows by applying $b_{1}=b_{2}=b_{3}=b$, Lemma 4.7, Proposition 4.8 and the symmetry of $[ijk]$ to the expression in (3.34). In Case 1, $b-2c_{i}=0$ is only possible if $i=3$, in which case $d_{1}=d_{2}=0$ while $d_{3}\neq 0$; in Case 2, only $b-2c_{2}$ can be zero, in which case $d_{1}=d_{3}=0$ while $d_{2}\neq 0$; and in Case 3, only $b-2c_{1}$ can be zero, in which case $d_{2}=d_{3}=0$ while $d_{1}\neq 0$. By Proposition 4.4, $(2,3)$ is $1$-selected, $(1,3)$ is $2$-selected, and $(1,2)$ is $3$-selected; see (3.35). We can thus use Proposition 3.26 to compute $r_{1}$ and $r_{2}$ in Case 1 if $2c_{3}=b$; $r_{1}$ and $r_{3}$ in Case 2 if $2c_{2}=b$; and $r_{2}$ and $r_{3}$ in Case 3 if $2c_{1}=b$. In every case, the desired expression for $r_{i}$ follows. Finally, the expression for $S$ follows by applying Lemma 4.7, Proposition 4.8 and the symmetry of $[ijk]$ to the expression in Proposition 3.25. ∎
@@ 4.4.3 Einstein metrics
[…]
@@
First suppose $c_{1}=0$. If $c_{2}+c_{3}=0$, then, in fact, $c_{2}=c_{3}=0$, so $c=0$ while $g$ is unconstrained. If $c_{2}+c_{3}\neq 0$ and $c_{2}c_{3}=0$, then $c=0$ and $g\in\{(x_{1},x_{2},x_{2}\pm x_{1})\mid x_{2}\neq\mp x_{1};\,x_{1}x_{2}\neq 0\}$. If $c_{2}+c_{3}\neq 0$ and $c_{2}c_{3}\neq 0$, then (S1) holds, that is, $c=\frac{1}{\lambda}$ and $g=\lambda(c_{2}+c_{3},c_{2},c_{3})$, $\lambda\in\mathbb{R}^{\times}$.
Now suppose $c_{2}+c_{3}=0$. Since $c_{1}=0$ was covered above, we also suppose $c_{1}\neq 0$. If $c_{3}=0$, then $c=0$ and $g\in\{(x_{1},x_{2},x_{1}-x_{2})\mid x_{2}\neq x_{1};\,x_{1}x_{2}\neq 0\}$. If $c_{3}\neq 0$, then (S4) holds, that is, $c=\frac{c_{1}^{2}}{4\lambda c_{3}^{2}}$ and $g=\lambda(c_{1},c_{2},c_{3})$, $\lambda\in\mathbb{R}^{\times}$, provides a solution. However, this is not the only solution if $c_{3}\notin\{0,c_{1}\}$, as then (S3) also holds, that is, $c=\frac{1}{\lambda}$ and $g=\lambda(c_{1},c_{2},c_{1}+c_{2})$, $\lambda\in\mathbb{R}^{\times}$, provides a solution. Further, if $c_{3}\notin\{0,c_{1},-c_{1}\}$, then (S2) also holds, that is, $c=\frac{1}{\lambda}$ and $g=\lambda(c_{1},c_{1}+c_{3},c_{3})$, $\lambda\in\mathbb{R}^{\times}$, provides a solution.

### v2

@@ 4.2 Compact real forms
To construct a flag supermanifold $M=(G_{0},\mathfrak{g})/(K_{0},\mathfrak{k})$ from a given circled Dynkin diagram, we first specify the compact real form $\mathfrak{g}$. In the following, we recall Dynkin diagrams, star operations and compact real forms for the complex Lie superalgebras of types $A$ and $C$. It would be interesting to study (e.g. using Vogan diagrams [Chu13]) how the construction changes for non-compact real forms.
@@
where $E_{ij}]=0$ if $i,j\leq m$ or $i,j>m$, and $E_{ij}]=1$ otherwise. Let $\mathfrak{h}^{\mathbb{C}}$ be the Cartan subalgebra spanned by all diagonal matrices, and denote by $\{\varepsilon_{1},\ldots,\varepsilon_{m+n}\}$ the $(\mathfrak{h}^{\mathbb{C}})^{\vee}$-basis where $\varepsilon_{i}(E_{jj})=\delta_{ij}$ for all $i,j$. For convenience, we introduce $\delta_{\mu}:=\varepsilon_{m+\mu}$ for each $\mu=1,\ldots,n$.
Let $\mathfrak{sl}(m|n)^{\mathbb{C}}:=\{X\in\mathfrak{gl}(m|n)^{\mathbb{C}}\,|\,\operatorname{Str}(X)=0\}$, and let $\mathfrak{b}^{\mathbb{C}}$ be the Borel subsuperalgebra spanned by all upper-triangular matrices. Then, the set of positive roots partitions as $\Delta^{+}=\Delta_{\bar{0}}^{+}\cup\Delta_{\bar{1}}^{+}$, where
$\Delta_{\bar{0}}^{+}=\{\varepsilon_{i}-\varepsilon_{j},\delta_{\mu}-\delta_{\nu}\,|\,1\leq i<j\leq m;1\leq\mu<\nu\leq n\},\qquad\Delta_{\bar{1}}^{+}=\{\varepsilon_{i}-\delta_{\mu}\,|\,1\leq i\leq m;1\leq\mu\leq n\},$
@@
$\Pi=\{\alpha_{i}=\varepsilon_{i}-\varepsilon_{i+1},\alpha_{m}=\varepsilon_{m}-\delta_{1},\alpha_{m+\mu}=\delta_{\mu}-\delta_{\mu+1}\,|\,1\leq i\leq m-1;1\leq\mu\leq n-1\}$
@@
The numbers in the Dynkin diagram indicate the multiplicities in the decomposition of the highest root: $\varepsilon_{1}-\delta_{n}=\alpha_{1}+\cdots+\alpha_{m+n-1}$. The Weyl vector follows from
@@
$\{\imath E_{ii}\,|\,i=1,\ldots,m+n\}\cup\{A_{ij},B_{ij}\,|\,1\leq i<j\leq m\}\cup\{A_{ij},B_{ij}\,|\,m+1\leq i<j\leq m+n\}$
$\cup\,\{\sqrt{\imath}A_{ij},\sqrt{\imath}B_{ij}\,|\,1\leq i\leq m;\,m+1\leq j\leq m+n\},$
@@
$\Delta$ $=\big\{\!\pm\!(\varepsilon_{1}-\varepsilon_{2}),\pm(\varepsilon_{2}-\varepsilon_{3}),\pm(\varepsilon_{1}-\varepsilon_{3}),\pm(\delta_{1}-\delta_{2})\big\}\cup\big\{\!\pm\!(\varepsilon_{i}-\delta_{1}),\pm(\varepsilon_{i}-\delta_{2})\,|\,i=1,2,3\big\},$
$\Delta_{K}$ $=\big\{\!\pm\!(\varepsilon_{1}-\varepsilon_{2}),\pm(\delta_{1}-\delta_{2})\big\}\cup\big\{\!\pm\!(\varepsilon_{3}-\delta_{1}),\pm(\varepsilon_{3}-\delta_{2})\big\},$
@@
$\Delta_{M}=\Delta-\Delta_{K}=\big\{\!\pm\!(\varepsilon_{i}-\varepsilon_{3}),\pm(\varepsilon_{i}-\delta_{1}),\pm(\varepsilon_{i}-\delta_{2})\,|\,i=1,2\big\},$
@@
We extend the analysis above to general $\operatorname{SU}(m|n)$ with a single circled node in the Dynkin diagram in Section 4.3, and with two circled nodes in Section $4.$ A similar analysis is carried out for $\operatorname{SOSp}(2|2n)$ in Section 4.5, although for a single circled node only.
@@
$\mathfrak{osp}(2|2n)^{\mathbb{C}}:=\big\{X\in\mathfrak{gl}(2|2n)^{\mathbb{C}}\,|\,\omega(Xu,v)+(-1)^{[X][u]}\omega(u,Xv)=0\text{ for all }u,v\in V\big\}.$
@@
$\mathfrak{h}^{\mathbb{C}}=\big\{\mathrm{diag}(a,-a,f_{1},\dots,f_{n},-f_{1},\dots,-f_{n})\,|\,a,f_{1},\ldots,f_{n}\in\mathbb{C}\big\}.$ (4.10)
@@
$\Delta_{\bar{0}}^{+}=\{\delta_{i}\pm\delta_{j}\,|\,1\leq i<j\leq n\}\cup\{2\delta_{j}\,|\,1\leq j\leq n\},\qquad\Delta_{\bar{1}}^{+}=\{\varepsilon\pm\delta_{j}\,|\,1\leq j\leq n\}.$
@@
[…]
@@ 4.4.1 Isotropy summands
$\mathfrak{m}_{k_{1},k_{2}}^{\mathbb{C}}:=\bigoplus_{\alpha\in\Delta_{k_{1},k_{2}}}\mathfrak{g}_{\alpha}^{\mathbb{C}},\qquad\Delta_{k_{1},k_{2}}:=\Big\{\sum_{i=1}^{m+n-1}c_{i}\alpha_{i}\in\Delta_{M}\,|\,c_{p}=k_{1},c_{q}=k_{2}\Big\},\qquad(k_{1},k_{2})\in I_{\mathfrak{m}^{\mathbb{C}}},$
@@
so the coefficient of $\alpha_{p}$ is $1$, while the coefficient of $\alpha_{q}$ is 0. This implies that $\mathfrak{m}^{{}^{\prime}\mathbb{C}}_{1,0}\subseteq\mathfrak{m}^{\mathbb{C}}_{1,0}$. Similarly, it follows that $\mathfrak{m}^{{}^{\prime}\mathbb{C}}_{k_{1},k_{2}}\subseteq\mathfrak{m}^{\mathbb{C}}_{k_{1},k_{2}}$ for all $(k_{1},k_{2})\in I_{\mathfrak{m}^{\mathbb{C}}}$, so
@@
As in (2.4), $\mathfrak{m}_{\bar{0}}$ is spanned by (1)(-1)-eigenvectors of the star operation restricted to $\mathfrak{m}_{\bar{0}}^{\mathbb{C}}$. Using Proposition 4.2, the star operation on $\mathfrak{m}_{\bar{0}}^{\mathbb{C}}$ splits into restrictions on $(\mathfrak{m}_{-1,0}^{\mathbb{C}}\oplus\mathfrak{m}_{1,0}^{\mathbb{C}})_{\bar{0}}$, $(\mathfrak{m}_{0,-1}^{\mathbb{C}}\oplus\mathfrak{m}_{0,1}^{\mathbb{C}})_{\bar{0}}$ and $(\mathfrak{m}_{-1,-1}^{\mathbb{C}}\oplus\mathfrak{m}_{1,1}^{\mathbb{C}})_{\bar{0}}$, respectively. For each pair $(k_{1},k_{2})\in\{(1,0),(0,1),(1,1)\}$, the real (1)(-1)-eigenspace of the star operation on $(\mathfrak{m}_{-k_{1},-k_{2}}^{\mathbb{C}}\oplus\mathfrak{m}_{k_{1},k_{2}}^{\mathbb{C}})_{\bar{0}}$ is spanned by $A_{\alpha}$ and $B_{\alpha}$, $\alpha\in(\Delta_{k_{1},k_{2}}^{+})_{\bar{0}}$, so
@@
$\{A_{\alpha},B_{\alpha}\,|\,\alpha\in\Delta_{k_{1},k_{2}}\cap\Delta_{\bar{0}}^{+}\}\cup\{\sqrt{\imath}A_{\alpha},\sqrt{\imath}B_{\alpha}\,|\,\alpha\in\Delta_{k_{1},k_{2}}\cap\Delta_{\bar{1}}^{+}\}$
@@ 4.4.2 Structure constants and Ricci curvature
Case 1: $p<q\leq m$ mnpm-n-p mn+pqm-n+p-q $q$
Case 2: $p<m<q$ mnpm-n-p mn+p+q-m-n+p+q 2mq2m-q
Case 3: $m\leq p<q$ mn+p-m-n+p mnp+qm-n-p+q 2mq2m-q
@@
If $123]\neq 0$, then $d_{1},d_{2},d_{3}$ are nonzero and the desired expression for $r_{i}$ follows by applying $b_{1}=b_{2}=b_{3}=b$, Lemma 4.7, Proposition 4.8 and the symmetry of $[ijk]$ to the expression in (3.34). In Case 1, $b-2c_{i}=0$ is only possible if $i=3$, in which case $d_{1}=d_{2}=0$ while $d_{3}\neq 0$; in Case 2, only $b-2c_{2}$ can be zero, in which case $d_{1}=d_{3}=0$ while $d_{2}\neq 0$; and in Case 3, only $b-2c_{1}$ can be zero, in which case $d_{2}=d_{3}=0$ while $d_{1}\neq 0$. By Proposition 4.4, $(2,3)$ is $1$-selected, $(1,3)$ is $2$-selected, and $(1,2)$ is $3$-selected; see (3.35). We can thus use Proposition 3.28 to compute $r_{1}$ and $r_{2}$ in Case 1 if $2c_{3}=b$; $r_{1}$ and $r_{3}$ in Case 2 if $2c_{2}=b$; and $r_{2}$ and $r_{3}$ in Case 3 if $2c_{1}=b$. In every case, the desired expression for $r_{i}$ follows. Finally, the expression for $S$ follows by applying Lemma 4.7, Proposition 4.8 and the symmetry of $[ijk]$ to the expression in Proposition 3.27. ∎
@@ 4.4.3 Einstein metrics
$(\mathrm{S1}):\quad(x_{1},x_{2},x_{3})=(c_{2}+c_{3},c_{2},c_{3}),\qquad c=1,$
$(\mathrm{S2}):\quad(x_{1},x_{2},x_{3})=(c_{1},c_{1}+c_{3},c_{3}),\qquad c=1,$
$(\mathrm{S3}):\quad(x_{1},x_{2},x_{3})=(c_{1},c_{2},c_{1}+c_{2}),\qquad c=1,$
$(\mathrm{S4}):\quad(x_{1},x_{2},x_{3})=(c_{1},c_{2},c_{3}),\qquad\qquad c=1+\frac{(b-2c_{1})(b-2c_{2})(b-2c_{3})}{4c_{1}c_{2}c_{3}}.$
@@
First suppose $c_{1}=0$. If $c_{2}+c_{3}=0$, then, in fact, $c_{2}=c_{3}=0$, so $c=0$ while $g$ is unconstrained. If $c_{2}+c_{3}\neq 0$ and $c_{2}c_{3}=0$, then $c=0$ and $g\in\{(x_{1},x_{2},x_{2}\pm x_{1})\,|\,x_{2}\neq\mp x_{1};\,x_{1}x_{2}\neq 0\}$. If $c_{2}+c_{3}\neq 0$ and $c_{2}c_{3}\neq 0$, then (S1) holds, that is, $c=\frac{1}{\lambda}$ and $g=\lambda(c_{2}+c_{3},c_{2},c_{3})$, $\lambda\in\mathbb{R}^{\times}$.
Now suppose $c_{2}+c_{3}=0$. Since $c_{1}=0$ was covered above, we also suppose $c_{1}\neq 0$. If $c_{3}=0$, then $c=0$ and $g\in\{(x_{1},x_{2},x_{1}-x_{2})\,|\,x_{2}\neq x_{1};\,x_{1}x_{2}\neq 0\}$. If $c_{3}\neq 0$, then (S4) holds, that is, $c=\frac{c_{1}^{2}}{4\lambda c_{3}^{2}}$ and $g=\lambda(c_{1},c_{2},c_{3})$, $\lambda\in\mathbb{R}^{\times}$, provides a solution. However, this is not the only solution if $c_{3}\notin\{0,c_{1}\}$, as then (S3) also holds, that is, $c=\frac{1}{\lambda}$ and $g=\lambda(c_{1},c_{2},c_{1}+c_{2})$, $\lambda\in\mathbb{R}^{\times}$, provides a solution. Further, if $c_{3}\notin\{0,c_{1},-c_{1}\}$, then (S2) also holds, that is, $c=\frac{1}{\lambda}$ and $g=\lambda(c_{1},c_{1}+c_{3},c_{3})$, $\lambda\in\mathbb{R}^{\times}$, provides a solution.